# On the Robustness of Kernel Goodness-of-Fit Tests

### v3

@@ Abstract
Goodness-of-fit testing is often criticized for its lack of practical relevance: since “all models are wrong”, the null hypothesis that the data conform to our model is ultimately always rejected as sample size grows. Despite this, probabilistic models are still used extensively, raising the more pertinent question of whether the model is good enough for the task at hand. This question can be formalized as a robust goodness-of-fit testing problem by asking whether the data were generated from a distribution that is a mild perturbation of the model. In this paper, we show that existing kernel goodness-of-fit tests are not robust under common notions of robustness including both qualitative and quantitative robustness. We further show that robustification techniques using tilted kernels, while effective in the parameter estimation literature, are not sufficient to ensure both types of robustness in the testing setting. To address this, we propose the first robust kernel goodness-of-fit test, which resolves this open problem by using kernel Stein discrepancy (KSD) balls. This framework encompasses many well-known perturbation models, such as Huber’s contamination and density-band models.
@@ 1 Introduction
This conundrum has led to the development of robust tests (Huber,, 1965; LeCam,, 1973; Dabak and Johnson,, 1994; Fauß et al.,, 2021), which aim to control the Type-I error rate (the probability of rejecting the null hypothesis) under a composite null hypothesis $H^{\mathrm{C}}_{0}:Q\in\mathcal{P}_{0}$, where $\mathcal{P}_{0}$ is a family of probability distributions, called the uncertainty set (Fauß et al.,, 2021). This uncertainty set is constructed to include $P$ and distributions close to $P$ in some appropriate notion of similarity. By designing $\mathcal{P}_{0}$ to include potential contaminations, such as adversarial contaminations or outliers, the composite null $H^{\mathrm{C}}_{0}$ can now hold in realistic scenarios, thus avoiding the aforementioned conundrum and aligning statistical and practical significance.
@@
Unfortunately, robust kernel GOF tests are absent from the existing literature on kernel tests. The closest work is that of Sun and Zou, (2023); Schrab and Kim, (2024), who proposed robust tests using uncertainty sets defined by the Maximum Mean Discrepancy (MMD; Müller, 1997; Gretton et al., 2012). These tests require samples from both $Q$ and $P$ and are thus applicable when simulating from $P$ is feasible (see, e.g., Lloyd and Ghahramani, 2015), but this would likely lead to significant additional computational cost and a lack of statistical efficiency due to the extra error induced by simulations.
We propose a novel class of robust kernel GOF tests based on the kernel Stein discrepancy (KSD). These tests are appropriate in the one-sample setting, where samples are available only from $Q$, and only an unnormalized density is available for $P$. To guarantee robustness against mild perturbations, we make use of tilted kernels proposed in a related construction for robust parameter estimation in Barp et al., (2019); Matsubara et al., (2022). We then study qualitative and quantitative robustness, two notions of robustness which capture a test’s insensitivity to perturbations in a neighborhood of $P$, as formally introduced in Section 2.3. Our contributions are summarized in Table 1 and detailed below:
@@
    In Section 3, we study the robustness properties of existing KSD tests. We show that KSD tests with stationary kernels are not necessarily robust under infinitesimal model deviation (which we call qualitative robustness), and never robust against fixed classes of misspecified models (called quantitative robustness). We then show that qualitative robustness can be guaranteed by using appropriately tilted kernels, but this is not sufficient to guarantee quantitative robustness.
@@
    In Section 4, we propose a kernel GOF test that controls the Type-I error rate against all distributions in a KSD-ball of radius $\theta>0$ centered at $P$. This test is straightforward to implement since it is identical to existing KSD tests up to a shift of the test statistic by $\theta$. We then make recommendation on how to select $\theta$ in practice to ensure robustness to common perturbations such as Huber’s contaminations (which includes outliers) or density-band contaminations.
@@
Existing KSD test $+$ stationary kernels Theorem 1
Existing KSD test $+$ tilted kernels Theorem 3
Novel KSD tests proposed in this paper Theorem 4
@@ 2 Background
Let $\mathcal{P}(\mathbb{R}^{d})$ denote the set of probability measures on $\mathbb{R}^{d}$. Given an integer $r>0$, $\mathcal{C}^{r}$ is the set of functions $f:\mathbb{R}^{d}\to\mathbb{R}$ that are $r$-times continuously differentiable. The gradient of $f\in\mathcal{C}^{1}$ is denoted $\nabla f(\mathbf{x})=(\partial_{1}f(\mathbf{x}),\ldots,\partial_{d}f(\mathbf{x}))^{\top}$. The set $\mathcal{C}_{b}^{1}$ contains functions in $\mathcal{C}^{1}$ that are bounded and with bounded gradient, and the set $\mathcal{C}_{b}^{2}$ contains $f\in\mathcal{C}_{b}^{1}$ for which $\sup_{\mathbf{x}\in\mathbb{R}^{d}}|\nabla^{\top}\nabla f(\mathbf{x})|<\infty$, where $\nabla^{\top}\nabla f(\mathbf{x})=\sum_{j=1}^{d}\partial_{j}\partial_{j}f(\mathbf{x})$. For a function $k:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$ with two arguments, $\nabla_{i}k(\mathbf{x},\mathbf{x}^{\prime})$ will be used to denote its gradient with respect to the $i$-th argument respectively for $i=1,2$, and we define $\nabla_{1}^{\top}\nabla_{2}k(\mathbf{x},\mathbf{x}^{\prime})=\sum_{j=1}^{d}\partial_{1j}\partial_{2j}k(\mathbf{x},\mathbf{x}^{\prime})$, where $\partial_{ij}k(\mathbf{x},\mathbf{x}^{\prime})=[\nabla_{i}k(\mathbf{x},\mathbf{x}^{\prime})]_{j}$ for $i=1,2$ and $j=1,\ldots,d$. The set $\mathcal{C}^{(1,1)}$ (resp., $\mathcal{C}^{(1,1)}_{b}$) denotes all functions $k:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$ with $(\mathbf{x},\mathbf{x}^{\prime})\mapsto\partial_{1j}\partial_{2j}k(\mathbf{x},\mathbf{x}^{\prime})$ continuous (resp., continuous and bounded) for all $j=1,\ldots,d$.
@@ 2.1 The Kernel Stein Discrepancy
where we have overloaded the notation by writing $D^{2}:(\mathbb{R}^{d})^{n}\to[0,\infty)$ as a test statistic. An alternative estimator is the U-statistic, which differs from (2) by summing only over disjoint pairs $i\neq j$. Although the U-statistic is unbiased, it can take negative values. Therefore, in this paper, we focus on V-statistics, leaving a discussion on how to extend our results to U-statistics to Appendix C.
@@ 2.2 Kernel Goodness-of-Fit Testing
where $\mathbf{W}\coloneqq(W_{1},\ldots,W_{n})\sim\textrm{Multinomial}(n;1/n,\ldots,1/n)$. This procedure, called Efron’s bootstrap or weighted bootstrap (Arcones and Gine,, 1992; Janssen,, 1994), assigns a multinomial weight to each observation, which mimics recomputing the test statistic using data re-sampled from $\mathbb{X}_{n}$ with replacement. Other bootstrap methods such as wild bootstraps (Leucht and Neumann,, 2013; Shao,, 2010) are also viable, particularly when the samples are potentially correlated (Chwialkowski et al.,, 2016), but we focus on the weighted bootstrap since they perform similarly for independent samples, as shown in Section B.8. The test threshold is selected as the $(1-\alpha)$-quantile of the distribution of $D^{2}_{\mathbf{W}}(\mathbb{X}_{n})$ conditional on $\mathbb{X}_{n}$, i.e.,
[…]
@@
[…]
@@ 2.3 Robustness for Goodness-of-Fit Testing
[…]
where $\epsilon\in[0,1]$ is the maximal contamination ratio, and the probability measure $R$ acts as arbitrary contamination. This model is appropriate when practitioners believe that $P$ accurately describes all but a small proportion of the data. Huber’s models have been studied in the context of robust GOF testing including Huber, (1965); Qin and Priebe, (2017), as well as in robust estimation (Hampel,, 1974; Huber and Ronchetti,, 2011). Special interests lie in the case when $R$ is restricted to point masses, i.e., $R=\delta_{\mathbf{z}}$ for some $\mathbf{z}\in\mathbb{R}^{d}$, where robustness is often called bias-robustness (Huber and Ronchetti,, 2011). Beyond Huber’s models, $\mathcal{P}_{0}$ can also be chosen as density-band models (Kassam,, 1981), which assume the density of the data-generating distribution is close to the model density up to a small error (see Section 4.2). Moreover, $\mathcal{P}_{0}$ can be set to a ball defined via a statistical divergence or metric, such as Hellinger distance (LeCam,, 1973), Wasserstein metric (Gao et al.,, 2018), and Maximum Mean Discrepancy (Sun and Zou,, 2023).
@@
[…]
Intuitively, this notion of robustness asks how sensitive the test outcome is under infinitesimally small model deviation. The shrinking-size condition ensures that the probability of rejection under any $Q\neq P$ does not trivially approach one due to the consistency of the test. By choosing the $\mathcal{P}_{0}^{n}$ in Definition 1 to be Prokhorov balls (Prokhorov,, 1956), we would recover the qualitative robustness introduced in Rieder, (1982, Definition 2.1), which parallels the conventional notion of qualitative robustness for estimators as argued in Rieder, (1982, Remark 2). However, distributions in Prokhorov balls do not have a simple form, posing challenges to the analysis. Our definition extends it to a general sequence of neighborhood. In Section 3, we will choose $\mathcal{P}_{0}^{n}$ to be Huber’s contamination models, which both significantly simplifies the analysis and encompasses a wide range of relevant scenarios. We will show that, within Huber’s neighborhood, the standard KSD test is not qualitatively robust with stationary kernels, but is qualitatively robust with appropriately tilted kernels.
Moreover, our definition of qualitative robustness is tied to a sequence of shrinking neighborhood $\{P_{0}^{n}\}_{n=1}^{\infty}$. Clearly, if a test is qualitatively robust to a sequence $\{P_{0}^{n}\}_{n=1}^{\infty}$, then it is also qualitatively robust to any sequences that decay faster. Therefore, loosely speaking, the rate of decay of $\{P_{0}^{n}\}_{n=1}^{\infty}$ characterizes the degree of qualitative robustness. In Section 3, we will explicitly derive the rate required for the standard KSD test to maintain qualitative robustness.
@@
In Section 4, we will choose $\mathcal{P}_{0}$ to be a KSD ball centered at $P$. We choose KSD balls over neighborhood defined by other statistical divergences because (i) this choice naturally lends itself to a GOF test that is both easy to implement and guarantees robustness at little extra computational cost, and (ii) it is easy to select the radius of such balls so that our proposed test is quantitatively robust to various types of contamination of interest, such as Huber’s contaminations or density-band contaminations.
@@ 3 The (Lack of) Robustness of Existing Kernel Goodness-of-fit Tests
We will now study the robustness of standard KSD GOF tests using stationary kernels in Section 3.1, then study their tilted counterparts which are popular in the parameter estimation literature in Section 3.2.
@@ 3.1 Existing KSD Tests with Stationary Kernels are not Qualitatively Robust
Our first result states that contamination of Huber’s type can considerably affect the probability of the standard KSD test rejecting the null hypothesis $H_{0}:Q=P$. Our result holds with the bootstrap threshold defined in (4), and all probabilities are taken over the randomness of both the sample and the bootstrap weights $\mathbf{W}\sim\textrm{Multinomial}(n;1/n,\ldots,1/n)$. The proof is in Section A.1.
@@
Theorem 1 immediately implies that the standard KSD test is not qualitatively robust to any sequence of Huber’s models that satisfies the rate condition $\epsilon_{n}=o(n)^{-1}$. At first glance, the lack of qualitative robustness might seem trivial given that the KSD test is consistent, meaning that the rejection probability under any $Q\neq P$ approaches one as $n\to\infty$. However, one subtlety is that the set $\mathcal{P}(P;\epsilon_{n})$ can also shrink in size, since $\epsilon_{n}$ is allowed to decay. The condition $\epsilon_{n}=o(n)^{-1}$ imposes a lower bound on this decay rate. Intuitively, this rate condition requires $\epsilon_{n}n$, the expected number of contamination in the sample, to grow with $n$. In particular, this excludes the case where $\epsilon_{n}n$ is a constant. Moreover, since Huber’s models $\mathcal{P}(P;\epsilon_{n})$ are contained within Prokhorov balls, Theorem 1 immediately implies non-qualitative robustness to Prokhorov neighborhoods.
@@
Theorem 1 is a consequence of the fact that the Stein kernel evaluated at $\mathbf{x}$, i.e., $\mathbf{x}\mapsto u_{p}(\mathbf{x},\mathbf{x})$, is unbounded when $P$ has an exploding score function. Exploding score functions are often associated with light tails. Examples of such distributions include those with a density of the form $p(\mathbf{x})\propto\exp(-\|\mathbf{x}\|_{2}^{r})$ with $r\geq 2$. These distributions have sub-Gaussian tails, and their score function has the form $\mathbf{s}_{p}(\mathbf{x})=-r\mathbf{x}\|\mathbf{x}\|_{2}^{r-2}$, which is unbounded for $r\geq 2$. On the other hand, this exploding-score condition excludes heavy-tailed models with bounded score functions, such as super-Laplacian or t-distributions (Gorham and Mackey,, 2017; Barp et al.,, 2024).
@@
The moment conditions in Theorem 1 are mild. For the example from Remark 1.1, namely $p(\mathbf{x})\propto\exp(-\|\mathbf{x}\|_{2}^{r})$ with $r\geq 2$, direct computation shows that these conditions hold if $h(\mathbf{x})$ and $\nabla h(\mathbf{x})$ decay at least as fast as $\||\mathbf{x}\|_{2}^{-(r-1)}$ so as to cancel the growth of the score $\|\mathbf{s}_{p}(\mathbf{x})\|_{2}=r\|\mathbf{x}\|_{2}^{r-1}$. Such kernels include squared-exponential kernels, IMQ kernels $h(u)=(1+\|u\|_{2}^{2})^{-s/2}$ with $s\geq r-1$, and Matérn kernels with sufficient smoothness. Section 5.1 will provide numerical evidence using a Gaussian model.
@@
[…]

### v4

@@ Abstract
Goodness-of-fit testing is often criticized for its lack of practical relevance: since “all models are wrong”, the null hypothesis that the data conform to our model is ultimately always rejected as the sample size grows. Despite this, probabilistic models are still used extensively, raising the more pertinent question of whether the model is good enough for the task at hand. This question can be formalized as a robust goodness-of-fit testing problem by asking whether the data were generated from a distribution that is a mild perturbation of the model. In this paper, we show that existing kernel goodness-of-fit tests are not robust under common notions of robustness including both qualitative and quantitative robustness. We further show that robustification techniques using tilted kernels, while effective in the parameter estimation literature, are not sufficient to ensure both types of robustness in the testing setting. To address this, we propose the first robust kernel goodness-of-fit test, which resolves this open problem by using kernel Stein discrepancy (KSD) balls. This framework encompasses many well-known perturbation models, such as Huber’s contamination and density-band models.
@@ 1 Introduction
This conundrum has led to the development of robust tests (Huber,, 1965; LeCam,, 1973; Dabak and Johnson,, 1994; Fauß et al.,, 2021), which aim to control the Type-I error rate (the probability of falsely rejecting the null hypothesis) under a composite null hypothesis $H^{\mathrm{C}}_{0}:Q\in\mathcal{P}_{0}$, where $\mathcal{P}_{0}$ is a family of probability distributions, called the uncertainty set (Fauß et al.,, 2021). This uncertainty set is constructed to include $P$ and distributions similar to $P$ in some appropriate notion. By designing $\mathcal{P}_{0}$ to include potential contaminations, such as adversarial contaminations or outliers, the composite null $H^{\mathrm{C}}_{0}$ can now hold in realistic scenarios, thus avoiding the aforementioned conundrum and aligning statistical and practical significance.
@@
Unfortunately, robust kernel GOF tests are absent from the existing literature on kernel tests. The closest work is that of Sun and Zou, (2023); Schrab and Kim, (2024), who proposed robust tests using uncertainty sets defined by the Maximum Mean Discrepancy (MMD; Müller, 1997; Gretton et al., 2012). These tests require samples from both $Q$ and $P$. However, sampling from $P$ might be infeasible or require significant additional computational cost. The approximation of $P$ by finite samples could also reduce statistical efficiency.
We propose a novel class of robust kernel GOF tests based on the kernel Stein discrepancy (KSD). These tests are appropriate in the one-sample setting, where samples are available only from $Q$, and an unnormalized density is available for $P$. To guarantee robustness against mild perturbations, we make use of tilted kernels proposed in a related construction for robust parameter estimation in Barp et al., (2019); Matsubara et al., (2022). We then study qualitative and quantitative robustness, two notions of robustness which capture a test’s insensitivity to perturbations in a neighborhood of $P$, as formally introduced in Section˜2.3. Our contributions are summarized in Table˜1 and detailed below:
@@
    In Section˜3, we study the robustness properties of existing KSD tests. We show that KSD tests with stationary kernels are not necessarily robust under infinitesimal model deviation (which we call qualitative robustness; cf. Definition˜1), and never robust against fixed classes of misspecified models (called quantitative robustness; cf. Definition˜2). We then show that qualitative robustness can be guaranteed by using appropriately tilted kernels, but this is not sufficient to guarantee quantitative robustness.
@@
    In Section˜4, we propose a kernel GOF test that controls the Type-I error rate against all distributions in a KSD-ball of radius $\theta>0$ centered at $P$. This test is straightforward to implement since it is identical to existing KSD tests up to a shift of the test statistic by $\theta$. We then make recommendation on how to select $\theta$ in practice to ensure robustness to common perturbations such as Huber’s contaminations (which includes outliers) or density-band contaminations.
@@
Existing KSD test ++ stationary kernels Theorem˜1
Existing KSD test ++ tilted kernels Theorem˜3
Novel KSD tests proposed in this paper Theorem˜4
@@ 2 Background
[…]
@@ 2.1 The Kernel Stein Discrepancy
where we have overloaded the notation by writing $D^{2}:(\mathbb{R}^{d})^{n}\to[0,\infty)$ as a test statistic. An alternative estimator is the U-statistic, which differs from (2) by summing only over disjoint pairs $i\neq j$. Although the U-statistic is unbiased, it can take negative values. Therefore, in this paper, we focus on V-statistics, leaving a discussion on how to extend our results to U-statistics to Appendix˜C.
@@ 2.2 Kernel Goodness-of-Fit Testing
where $\mathbf{W}\coloneqq(W_{1},\ldots,W_{n})\sim\textrm{Multinomial}(n;1/n,\ldots,1/n)$. This procedure, called Efron’s bootstrap or weighted bootstrap (Arcones and Gine,, 1992; Janssen,, 1994), assigns a multinomial weight to each observation, which mimics recomputing the test statistic using data re-sampled from $\mathbb{X}_{n}$ with replacement. Other bootstrap methods such as wild bootstraps (Leucht and Neumann,, 2013; Shao,, 2010) are also viable, particularly when the samples are potentially correlated (Chwialkowski et al.,, 2016), but we focus on the weighted bootstrap since they perform similarly for independent samples, as shown in Section˜B.8. The test threshold is selected as the $(1-\alpha)$-quantile of the distribution of $D^{2}_{\mathbf{W}}(\mathbb{X}_{n})$ conditional on $\mathbb{X}_{n}$, i.e.,
$q^{2}_{\infty,1-\alpha}(\mathbb{X}_{n})\;\coloneqq\;\inf\big{\{}u\in\mathbb{R}:\;1-\alpha\leq\Pr\nolimits_{\mathbf{W}}\big{(}D^{2}_{\mathbf{W}}(\mathbb{X}_{n})\leq u\;|\;\mathbb{X}_{n}\big{)}\big{\}}\;.$ (4)
@@
$q^{2}_{B,1-\alpha}(\mathbb{X}_{n})\;\coloneqq\;\inf\Bigg{\{}u\in\mathbb{R}:\;1-\alpha\leq\frac{1}{B+1}\bigg{(}\mathbbm{1}\{D^{2}(\mathbb{X}_{n})\leq u\}+\sum_{b=1}^{B}\mathbbm{1}\{D^{2}_{\mathbf{W}^{b}}(\mathbb{X}_{n})\leq u\}\bigg{)}\Bigg{\}}\;,$ (5)
@@ 2.3 Robustness for Goodness-of-Fit Testing
$\mathcal{P}(P;\epsilon)\;\coloneqq\;\{(1-\epsilon^{\prime})P+\epsilon^{\prime}R:\;0\leq\epsilon^{\prime}\leq\epsilon,\;R\in\mathcal{P}(\mathbb{R}^{d})\}\;,$ (6)
where $\epsilon\in[0,1]$ is the maximal contamination ratio, and the probability measure $R$ acts as arbitrary contamination. This model is appropriate when practitioners believe that $P$ accurately describes all but a small proportion of the data. Huber’s models have been studied in the context of robust GOF testing including Huber, (1965); Qin and Priebe, (2017), as well as in robust estimation (Hampel,, 1974; Huber and Ronchetti,, 2011). Special interests lie in the case when $R$ is restricted to point masses, i.e., $R=\delta_{\mathbf{z}}$ for some $\mathbf{z}\in\mathbb{R}^{d}$, where robustness is often called bias-robustness (Huber and Ronchetti,, 2011). Beyond Huber’s models, $\mathcal{P}_{0}$ can also be chosen as density-band models (Kassam,, 1981), which assume the density of the data-generating distribution is close to the model density up to a small error (see Section˜4.2). Moreover, $\mathcal{P}_{0}$ can be set to a ball defined via a statistical divergence or metric, such as Hellinger distance (LeCam,, 1973), Wasserstein metric (Gao et al.,, 2018), and Maximum Mean Discrepancy (Sun and Zou,, 2023).
@@
$\sup\nolimits_{Q\in\mathcal{P}_{0}^{n}}\big{|}\Pr\nolimits_{\mathbb{X}_{n}\sim Q}\big{(}T_{n}(\mathbb{X}_{n})>\gamma_{n}(\mathbb{X}_{n})\big{)}-\Pr\nolimits_{\mathbb{X}_{n}^{\ast}\sim P}\big{(}T_{n}(\mathbb{X}_{n}^{\ast})>\gamma_{n}(\mathbb{X}_{n}^{\ast})\big{)}\big{|}\;\to\;0\;.$ (7)
Intuitively, this notion of robustness asks how sensitive the test outcome is under infinitesimally small model deviation. The shrinking-size condition ensures that the probability of rejection under any $Q\neq P$ does not trivially approach one due to the consistency of the test. By choosing the $\mathcal{P}_{0}^{n}$ in Definition˜1 to be Prokhorov balls (Prokhorov,, 1956), we would recover the qualitative robustness introduced in Rieder, (1982, Definition 2.1), which parallels the conventional notion of qualitative robustness for estimators as argued in Rieder, (1982, Remark 2). However, distributions in Prokhorov balls do not have a simple form, posing challenges to the analysis. Our definition extends it to a general sequence of neighborhood. In Section˜3, we will choose $\mathcal{P}_{0}^{n}$ to be Huber’s contamination models, which both significantly simplifies the analysis and encompasses a wide range of relevant scenarios. We will show that, within Huber’s neighborhood, the standard KSD test is not qualitatively robust with stationary kernels, but is qualitatively robust with appropriately tilted kernels.
Moreover, our definition of qualitative robustness is tied to a sequence of shrinking neighborhood $\{P_{0}^{n}\}_{n=1}^{\infty}$. Clearly, if a test is qualitatively robust to a sequence $\{P_{0}^{n}\}_{n=1}^{\infty}$, then it is also qualitatively robust to any sequences that decay faster. Therefore, loosely speaking, the rate of decay of $\{P_{0}^{n}\}_{n=1}^{\infty}$ characterizes the degree of qualitative robustness. In Section˜3, we will explicitly derive the rate required for the standard KSD test to maintain qualitative robustness.
@@
In Section˜4, we will choose $\mathcal{P}_{0}$ to be a KSD ball centered at $P$. We choose KSD balls over neighborhood defined by other statistical divergences because (i) this choice naturally lends itself to a GOF test that is both easy to implement and guarantees robustness at little extra computational cost, and (ii) it is easy to select the radius of such balls so that our proposed test is quantitatively robust to various types of contamination of interest, such as Huber’s contaminations or density-band contaminations.
@@ 3 The (Lack of) Robustness of Existing Kernel Goodness-of-fit Tests
We will now study the robustness of standard KSD GOF tests using stationary kernels in Section˜3.1, then study their tilted counterparts which are popular in the parameter estimation literature in Section˜3.2.
@@ 3.1 Existing KSD Tests with Stationary Kernels are not Qualitatively Robust
Our first result states that contamination of Huber’s type can considerably affect the probability of the standard KSD test rejecting the null hypothesis $H_{0}:Q=P$. Our result holds with the bootstrap threshold defined in (4), and all probabilities are taken over the randomness of both the sample and the bootstrap weights $\mathbf{W}\sim\textrm{Multinomial}(n;1/n,\ldots,1/n)$. The proof is in Section˜A.1.
@@
Theorem˜1 immediately implies that the standard KSD test is not qualitatively robust to any sequence of Huber’s models that satisfies the rate condition $\epsilon_{n}=o(n)^{-1}$. At first glance, the lack of qualitative robustness might seem trivial given that the KSD test is consistent, meaning that the rejection probability under any $Q\neq P$ approaches one as $n\to\infty$. However, one subtlety is that the set $\mathcal{P}(P;\epsilon_{n})$ can also shrink in size, since $\epsilon_{n}$ is allowed to decay. The condition $\epsilon_{n}=o(n)^{-1}$ imposes a lower bound on this decay rate. Intuitively, this rate condition requires $\epsilon_{n}n$, the expected number of contamination in the sample, to grow with $n$. In particular, this excludes the case where $\epsilon_{n}n$ is a constant. Moreover, since Huber’s models $\mathcal{P}(P;\epsilon_{n})$ are contained within Prokhorov balls, Theorem˜1 immediately implies non-qualitative robustness to Prokhorov neighborhoods.
@@
Theorem˜1 is a consequence of the fact that the Stein kernel evaluated at $\mathbf{x}$, i.e., $\mathbf{x}\mapsto u_{p}(\mathbf{x},\mathbf{x})$, is unbounded when $P$ has an exploding score function. Exploding score functions are often associated with light tails. Examples of such distributions include those with a density of the form $p(\mathbf{x})\propto\exp(-\|\mathbf{x}\|_{2}^{r})$ with $r\geq 2$. These distributions have sub-Gaussian tails, and their score function has the form $\mathbf{s}_{p}(\mathbf{x})=-r\mathbf{x}\|\mathbf{x}\|_{2}^{r-2}$, which is unbounded for $r\geq 2$. On the other hand, this exploding-score condition excludes heavy-tailed models with bounded score functions, such as super-Laplacian or t-distributions (Gorham and Mackey,, 2017; Barp et al.,, 2024).
@@
The moment conditions in Theorem˜1 are mild. For the example from Remark˜1.1, namely $p(\mathbf{x})\propto\exp(-\|\mathbf{x}\|_{2}^{r})$ with $r\geq 2$, direct computation shows that these conditions hold if $h(\mathbf{x})$ and $\nabla h(\mathbf{x})$ decay at least as fast as $\||\mathbf{x}\|_{2}^{-(r-1)}$ so as to cancel the growth of the score $\|\mathbf{s}_{p}(\mathbf{x})\|_{2}=r\|\mathbf{x}\|_{2}^{r-1}$. Such kernels include squared-exponential kernels, IMQ kernels $h(u)=(1+\|u\|_{2}^{2})^{-s/2}$ with $s\geq r-1$, and Matérn kernels with sufficient smoothness. Section˜5.1 will provide numerical evidence using a Gaussian model.
@@
Let $\mathcal{Q}_{\delta}\coloneqq\{Q\in\mathcal{P}(\mathbb{R}^{d}):S(Q,P)\geq\delta\}$, where $S$ is some statistical divergence or metric. The separation boundary (Ingster,, 1987, 1993) of a test is the fastest decaying sequence $\{\delta_{n}\}_{n}$ such that the test power under any $Q\in\mathcal{Q}_{\delta_{n}}$ still converges to 1 as $n\to\infty$. From this perspective, Theorem˜1 implies that the separation boundary of the standard KSD test must decay at least with rate $\epsilon_{n}=o(n)^{-1}$, whenever $Q_{\delta_{n}}$ contains Huber-contamination models. This complements existing results on the separation boundary of KSD tests. The most relevant works are Schrab et al., (2022); Hagrass et al., (2025), where they consider alternative distributions $Q$ that have a density with respect to either the Lebesgue measure or the target model $P$. In particular, their results cannot be used to derive Theorem˜1, since their density assumption on $Q$ is violated in our case. This is because we consider alternatives of the form $Q=(1-\epsilon_{n})P+\epsilon_{n}\delta_{\mathbf{z}}$, which involve a Dirac delta measure and thus has no density with respect to the Lebesgue measure nor any continuous measures, including $P$.